OpenStudy (anonymous):

algebra 1 please help!!!

OpenStudy (anonymous):

\[\frac{4^4}{4^6}=\frac{1}{4^{6-4}}=\frac{1}{4^2}\]

OpenStudy (anonymous):

yes it does

OpenStudy (anonymous):

no for #4 you recall that \(b^0=1\)

OpenStudy (anonymous):

so then it is 5 to the 1st power, o the answer is 5?

OpenStudy (anonymous):

Or do it: \[\frac{ 5^{1} }{ 5^{1} }=5/5=1\]

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

\[5^{0}=\frac{5^{1}}{5^{1}}=\frac{5}{5}=1\]

OpenStudy (anonymous):

thanks!

OpenStudy (mertsj):

yw